OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? −x + 2y = 9 4x + 5y = 9 −9 + 2y −9 − 2y 9 − 2y 9 + 2y

jhonyy9 (jhonyy9):

how do you think ,any idea ,opinion ?

OpenStudy (anonymous):

no i have no idea :(

jhonyy9 (jhonyy9):

so there is -x +2y =9 yes ? on the first step i think that you can assuming to both sides x and check it please than what will get ?

jhonyy9 (jhonyy9):

please you cooperate anything ,ok ?

OpenStudy (anonymous):

ok ill try

jhonyy9 (jhonyy9):

so than write it here

jimthompson5910 (jim_thompson5910):

solve −x + 2y = 9 for x

jhonyy9 (jhonyy9):

-x +x +2y=9 +x

OpenStudy (anonymous):

i think its −9 − 2y @jhonyy9

jhonyy9 (jhonyy9):

not is right

jhonyy9 (jhonyy9):

i like learn to you step by step

OpenStudy (anonymous):

what? sorry i didnt understand

jimthompson5910 (jim_thompson5910):

−x + 2y = 9 −x = 9 - 2y ... Subtract 2y from both sides. x = -9 + 2y ... change the sign of every term to fully isolate x

OpenStudy (anonymous):

9 − 2y @jim_thompson5910 i think its this on

jimthompson5910 (jim_thompson5910):

no it's the one i wrote above

jimthompson5910 (jim_thompson5910):

9-2y is NOT the same as -9 + 2y

OpenStudy (anonymous):

-9 + 2y so it this one? :D

OpenStudy (anonymous):

@jim_thompson5910

jhonyy9 (jhonyy9):

yes

OpenStudy (anonymous):

thank you :)

jhonyy9 (jhonyy9):

yw